Suppose that an insurance agent offers you a policy that will provide you with a yearly income of $30,000 in 30 years. What is the comparable salary today, assuming an inflation rate of 6% compounded annually? (Round your answer to the nearest cent.)

Answer by rfer(16322)

P=30000(1.06)^-30
P=30000*0.17411
P=5223.30
----------------
30000=P(1.06)^30
30000/5.7434912=P
5223.30=P